### v1

# Dialogues with AI Reduce Beliefs in Misinformation but Build No Lasting Discernment Skills

Abstract.
Given the growing prevalence of fake information, including increasingly realistic AI-generated news, there is an urgent need to train people to better evaluate and detect misinformation. While interactions with AI have been shown to durably reduce people’s beliefs in false information, it is unclear whether these interactions also teach people the skills to discern false information themselves. We conducted a month-long study where 67 participants classified news headline-image pairs as real or fake, discussed their assessments with an AI system, followed by an unassisted evaluation of unseen news items to measure accuracy before, during, and after AI assistance. While AI assistance produced immediate improvements during AI-assisted sessions (+21% average), participants’ unassisted performance on new items declined significantly by week 4 (-15.3%). These results indicate that while AI may help immediately, it ultimately degrades long-term misinformation detection abilities.
1. Introduction
AI Assistants such as ChatGPT, Claude and Grok are increasingly used to evaluate the credibility of online information, from judging the authenticity of news headlines and viral images to answering whether medical claims (Shekar et al., 2024) or political rumors (Kuznetsova et al., 2025) are true. While recent research suggests such systems can reduce belief in specific false claims (Costello et al., 2024), it remains unclear whether these conversations teach humans to detect misinformation or merely shift beliefs about false information with AI assistance.
AI-generated misinformation is spreading rapidly online, often gaining widespread traction and causing significant societal harm. For example, AI-generated reports about an explosion near the Pentagon in May 2023 garnered thousands of shares before being debunked (See Appendix section A, Figure 9). This is mostly due to the cognitive biases making people susceptible to false but compelling narratives (Pennycook and Rand, 2019b).
While both human and machine-based fact-checking methods (Chuai et al., 2024a; Ecker et al., 2010) have been used to annotate and remove false information from online spheres, they struggle with the speed at which misinformation is produced and spread, or fail to be accurate. Additionally, machine learning-based detection algorithms (Wang et al., 2020, 2023; Sha et al., 2023) are computationally expensive and struggle to keep up with the speed at which misinformation can be generated from generative models. Humans are poor at detecting AI-generated content due to the increasing sophistication of generative models and a lack of knowledge about their outputs, such as inconsistent lighting or unnatural textures (Boutadjine et al., 2025). Adding to this challenge, research shows that even when false information is explicitly corrected with human or machine-based corrections (Chuai et al., 2024a; Ecker et al., 2010; Wang et al., 2020, 2023; Sha et al., 2023), people often retain their initial beliefs—a persistent cognitive bias known as the continual influence effect (Johnson and Seifert, 1994a).
While fact-checking systems have shown mixed results on helping people detect misinformation and correct their beliefs about it (Chuai et al., 2024b; Walter et al., 2020), recent work on AI chatbots has demonstrated that dialogues with an AI chatbot can reduce belief in conspiracy theories by 22% (Costello et al., 2024), suggesting potential for AI-assisted belief correction. Furthermore, AI chatbots for education have shown promise in various domains (Cai et al., 2024; Sabnis and Nagashima, 2024; Ceha et al., 2021; Chhibber and Law, 2019), particularly when acting as collaborative partners rather than authoritative sources. This makes the AI chatbot a potential tool to teach people how to detect and update their beliefs about misinformation. Yet emerging research on Human-AI interaction suggests that AI chatbots can foster overreliance (Danry et al., 2023; Kosmyna et al., 2025), with users deferring to system outputs even when incorrect (Danry et al., 2023), potentially undermining the development of independent reasoning skills (Buçinca et al., 2021; Gerlich, 2025). This raises an important question: Do dialogues with AI chatbots actually build lasting misinformation detection skills, or do they create dependency that undermines humans’ independent judgment?
We investigate this question by designing an AI chatbot prompted to generate persuasive dialogues for misinformation detection and by conducting a longitudinal study measuring users’ ability to detect misinformation both with and without AI assistance. We summarize our contributions as follows:
System Design: We design an AI chatbot that uses persuasive dialogue to help people evaluate news authenticity and detect misinformation. Figure 2 illustrates our system’s user interface.
Longitudinal Study: We conduct the first month-long, three-phase longitudinal study measuring misinformation detection effects both with and without AI assistance (N=67).
Dependency Paradox: We demonstrate a critical limitation where AI provides immediate accuracy gains (+21%) but leads to a significant decline in unassisted performance (-15.3% by week 4), particularly for detecting fake content.
Conversation Analysis: We perform a comprehensive analysis of the recorded conversations to automatically classify 24 conversational parameters, examining Human-AI interaction behaviors, providing insights into their correlation with learning outcomes versus dependency development.
Open Source System and Dataset: We will publicly release our AI chatbot, 7,203 Human-AI conversation pairs, and 4,536 participant judgments on news authenticity (Real or Fake and a confidence rating) to enable future research.
2. Related Work
2.1. Why Misinformation Persists
The proliferation of AI-generated news media content presents major threats to the reliability of online information ecosystems (Farouk and Fahmi, 2024; Zhang and Ghorbani, 2020). As AI generation tools become more sophisticated and accessible, the creation and dissemination of manipulated images alongside misleading headlines has increased dramatically across social media platforms (Giansiracusa, 2021). The combination of visual and textual misinformation content is particularly persuasive, as research has shown that humans tend to place heightened trust in visual evidence while often lacking the skills to detect subtle manipulations or critically evaluate accompanying headlines (Jagadish and Jasmine, 2024; Pennycook and Rand, 2019a; Vosoughi et al., 2018).
Decades of psychological research explain why misinformation is so resistant to correction once it is initially believed. Classic work on rumor and memory demonstrated that false claims continue to shape reasoning even after being explicitly retracted (Allport and Postman, 1947; Johnson and Seifert, 1994b), a phenomenon now termed the continued influence effect. Meta-analyses and reviews confirm that even detailed corrections rarely eliminate lingering effects (Lewandowsky et al., 2012; Ecker et al., 2010; Walter and Tukachinsky, 2020; Ecker et al., 2022).
Another well-documented mechanism is the illusory truth effect, where repetition increases processing fluency and thus perceived accuracy (Pennycook et al., 2018; Hassan and Barber, 2021; Fazio et al., 2015). Importantly, even prior knowledge does not reliably inoculate against this effect — people with strong factual knowledge are still vulnerable to repeated falsehoods (Fazio et al., 2015). Recent work shows that reminders and retractions can sometimes backfire, particularly when cognitive load is high or attention is divided (Sanderson et al., 2023). To overcome the continued influence effect, simply correcting false information is insufficient. Research demonstrates that people need coherent alternative explanations that fill the explanatory void left behind to effectively remove the influence of misinformation (Ecker et al., 2010; Johnson and Seifert, 1994b; Rich and Zaragoza, 2016; Ithisuphalap et al., 2020).
Together, these findings highlight why misinformation proves so resistant to correction. Heuristics and intuitive reasoning further amplify this vulnerability, as people often rely on fast, effortless judgments rather than deliberate scrutiny (Tversky and Kahneman, 1974; Thompson et al., 2011; Kahneman, 2011). This persistence underscores the need for interventions that not only correct specific falsehoods but also address the psychological mechanisms that sustain them.
At the same time, psychological evidence indicates that people can be trained to think more critically and resist misinformation. Instruction in informal fallacies improves fake news detection (Hruschka and Appel, 2023; Neuman and Weizman, 2003), while dispositions toward actively open-minded thinking are associated with reduced susceptibility (Stanovich and Toplak, 2023). Encouraging deliberation also helps: prompting individuals to reflect systematically lowers belief in false news headlines (Bago et al., 2020). These findings suggest that some resilience against misinformation is achievable.
2.2. Interventions to Reduce Misinformation Beliefs
Extensive research has examined how to mitigate misinformation persistence through corrective information provided before, during, or after exposure. Professional fact-checks and platform warning labels have been found to reduce belief in false stories, but often to a modest effect (Chuai et al., 2024a). Even when corrections are well-designed, people frequently continue to draw inferences from the original misinformation, a phenomenon captured by the continued influence effect (Ecker et al., 2010; Rich and Zaragoza, 2016; Walter and Tukachinsky, 2020). Meta-analyses confirm that while fact-checks do exert a positive influence, their success depends heavily on factors such as timing, repetition, context, and cognitive load during exposure (Ecker et al., 2017; Ithisuphalap et al., 2020; Sanderson et al., 2023).
Recognizing the limitations of post-hoc correction, researchers have explored more proactive strategies. Inoculation or “prebunking” approaches aim to build resilience by warning people about misleading techniques or exposing them to weakened doses of misinformation in advance. This method has shown promise in buffering individuals against subsequent falsehoods (Lewandowsky et al., 2012; Van der Linden et al., 2017; Pennycook and Rand, 2021), but it requires anticipating specific manipulative tactics and often struggles to generalize across the wide variety of misinformation domains.
A complementary line of work targets decision heuristics at the moment of judgment. Subtle prompts that draw attention to accuracy increase discernment both when evaluating and when sharing content online (Pennycook et al., 2020, 2021; Pennycook and Rand, 2022; Danry et al., 2020). Similarly, simple accuracy nudges have been found to reduce the spread of false information on social media at scale. Yet these interventions, like corrections and prebunking, typically operate in a one-shot, non-interactive manner. They provide external inputs but rarely engage with how individuals themselves reason about a claim.
2.3. AI Dialogues and Persuasion
Dialogue-based interventions differ from static fact-checks by responding to the specific reasons people give for their beliefs. In the strongest demonstration to date, Costello et al. asked participants to describe why they endorsed a conspiracy claim and then instructed the AI chatbot to reply persuasively, but respectfully, with counterarguments tailored to those reasons (Costello et al., 2024). The result was a large and durable reduction in conspiracy belief. Follow-up analyses clarify the mechanism: the dialogues were effective when the AI chatbot provided factual counterevidence and alternative explanations targeted to the user’s arguments; when the system was prompted to persuade without presenting counterevidence, the effect disappeared (Costello et al., [n. d.]). In short, the intervention worked not because an AI was present, but because users experienced having their own arguments acknowledged and then met with “just the facts” that directly undermined those arguments.
Evidence from adjacent AI chatbot system settings is consistent with this evidence-centric account. Studies of AI news chatbots show that perceived trust and persuasiveness increase when the agent confidently presents evidence-backed claims, shaping users’ perceptions of sources and narratives (Govers et al., 2025). Work on conversational mediation similarly finds that prompt-tuned dialogue strategies can steer how people evaluate arguments during contentious debates (Govers et al., 2024). Across these lines of work, the common thread is a persuasive conversational stance of the AI chatbots that supplies clear, tailored reasons and sources to move users toward accurate conclusions. While this work shows that conversations with AI chatbots are capable of changing people’s beliefs to make them believe misinformation less, it is unclear how much of these effects are caused by reliance on the AI chatbot system or the user internalizing the skills to make good discernments independently.
2.4. Overreliance on AI and the Question of Learning
While persuasive, fact-focused dialogues can correct beliefs in the moment, there is a growing concern that such answer-giving strategies foster reliance on the AI rather than cultivating independent reasoning. Evidence from across HCI and psychology points to this risk. Studies of AI-assisted decision-making repeatedly find that users defer to system outputs even when they are erroneous, a phenomenon termed overreliance (Buçinca et al., 2021; Poursabzi-Sangdeh et al., 2021). Providing explanations or rationales does not consistently mitigate this effect; in some cases, explanations increase deference by making outputs appear more authoritative (Poursabzi-Sangdeh et al., 2021).
Recent work in Human–AI interaction illustrates the same dynamic in high-stakes reasoning contexts. Participants using AI systems that confidently “tell” rather than “ask”, report going along with the system because it sounds knowledgeable, thereby reducing their own engagement in critical evaluation (Danry et al., 2023). Similar patterns are observed in writing and analysis tasks: generative AI tools homogenize style (Agarwal et al., 2025), reduce self-reported cognitive effort (Lee et al., 2025), and accumulate what has been called cognitive debt, where reliance grows and effort declines over repeated use (Kosmyna et al., 2025). These observations align with classic automation-bias and algorithm-(aversion/appreciation) results showing systematic deference to automated recommendations under authority cues and cognitive load (Parasuraman and Riley, 1997; Dietvorst et al., 2015; Logg et al., 2019), and with cautionary findings from wearable guidance that directive cues invite default-to-system behavior unless prompts are designed to elicit reflection (Danry et al., 2020, 2023).
Taken together, these findings suggest that the very strategies that make persuasive dialogues effective—providing confident, tailored factual answers—might also be those most likely to induce reliance and discourage the type of cognitive engagement that might lead to learning (Buçinca et al., 2021). In the misinformation domain, this raises a critical open question: do AI dialogues merely persuade users about specific claims in the moment, or do they help users develop misinformation detection skills to evaluate new claims on their own? Our study directly addresses this gap by separating immediate gains during assisted interactions from unaided performance on novel misinformation, testing whether persuasive AI dialogues foster discernment or dependency.
3. System Design
We designed a web-based AI chatbot to investigate whether dialogues with AI can help humans learn to distinguish between real and fake news content over the course of a month-long study. The system presents participants with real and AI-generated news headline-image pairs, then engages them in conversations designed to guide them toward accurate authenticity assessment. In the following, we discuss the system overview, data curation, and system evaluation.
3.1. System Overview
To evaluate the effects of an AI chatbot on misinformation discernment, our system implementation closely follows the state-of-the-art approach (Costello et al., 2024) for generating persuasive dialogues. Combining LLMs with web search has proven to improve models’ accuracy for fact-related queries (Nakano et al., 2021; Alzubi et al., 2025). We therefore integrated web search for headlines and artifact detection for images, adapting the prompt from Costello et al. (2024) to improve our system’s accuracy on current misinformation examples in our dataset. The system consists of three main components: (i) React-based front-end for user interface, (ii) OpenAI GPT-4o prompted for persuasive dialogue generation (prompts can be found in Appendix B), and (iii) Google Custom Search API for real-time information retrieval. Figure 3 illustrates the system architecture. The React front-end manages the user interface, presenting news headline-image pairs and collecting participant responses, including initial authenticity ratings and confidence levels (See Figure 2). Those ratings are then sent to an AI chatbot system running GPT-4o, prompted to generate contextually appropriate persuasive responses to persuade users to change their beliefs about the headline-image pair if they were wrong (refer to Figure 13 to view an example from one such conversation). If they were correct, the AI response reinforces the user’s belief, and the user can still interact with the AI system (refer to Figure 12 to view an example from one such conversation).
This is implemented through a structured conversational approach where GPT-4o: (1) acknowledges the participant’s reasoning, (2) presents relevant factual evidence retrieved from Google Search, and (3) guides users through iterative persuasive dialogue toward more accurate assessments of the news item’s authenticity. The system architecture supports up to 9 rounds of conversation per news item. The conversation stops when 9 rounds are completed, or when the AI analyzes from the conversation that the user is done talking about the current news item, and then it moves to the next news item. All interactions are logged, including participant responses, AI-generated content, and metadata for subsequent analysis. Data is stored using both Qualtrics integration and Google Sheets API for analysis.
3.2. Data Curation
We randomly selected 55 news items from the MiRAGeNews dataset (Huang et al., 2024), which comprises 12,500 high-quality image-caption pairs—both real news items and AI-generated pairs produced using state-of-the-art generators—along with their ground truth labels (Real or Fake). Figure 4 shows representative examples from the dataset. These selected news items were processed through the architecture discussed in Figure 3 and prompted to output ”Real” or ”Fake”. The output was then compared with the ground truth labels from the MiRAGeNews dataset. The system misclassified 6 items, which we removed from the dataset. The remaining 49 items formed our final dataset, which was distributed across three study phases. Throughout the three phases, participants evaluated 4 randomly selected news items (2 real and 2 fabricated), followed by confidence ratings on 4 previously unseen items to assess whether they could correctly distinguish between real and fake news in unassisted sessions. In phases 2 and 3, the unassisted evaluation sessions included 6 items: 4 novel unseen items and 2 items previously encountered in the preceding phase. This design allowed us to measure both immediate learning on misinformation discernment and retention over time. The items span diverse topics, including elections, protests, court proceedings, diplomatic events, health matters, and disasters, ensuring broad coverage of typical news categories where misinformation commonly appears.
3.3. System Evaluation
Before deploying the system with participants, we conducted extensive validation of the AI’s accuracy to ensure reliable performance. We initially tested our system on a sample of eight items across three different prompt variations using GPT-4o under four different conditions based on available information: (i) Google search on headline with images passed to GPT-4o, (ii) Google search on headline without images, (iii) no Google search with images on GPT-4o, and (iv) no Google search and headline only on GPT-4o. We tested prompts designed for artifact detection only, persuasion/belief change only, and a combined forensic expert and persuasion approach. Figure 10 contains the detailed prompts, and Table 1 contains the evaluation metrics for all conditions.
| Prompt | Evaluation Metrics |
| Prompt 1: News Forensic Expert (Artifact detection only) | Overall Accuracy | 71.43% | 78.57% | 35.71% | 0.00% | ||||||||||||
| Rejection Rate | 21.43% | 21.43% | 50.00% | 92.86% | |||||||||||||
| Accuracy (Non Rejected) | 90.91% | 100.00% | 71.43% | 0.00% | |||||||||||||
| Prompt 2: Persuasion (Persuasive Prompting) | Overall Accuracy | 71.40% | 78.57% | 28.57% | 0.00% | ||||||||||||
| Rejection Rate | 14.20% | 14.20% | 78.57% | 92.00% | |||||||||||||
| Accuracy (Non Rejected) | 83.30% | 91.60% | 66.66% | 0.00% | |||||||||||||
| Prompt 3: Combined (News Forensic Expert + Persuasion) | Overall Accuracy | 100% | 78.5% | 42.8% | 0% | ||||||||||||
| Rejection Rate | 0% | 14.20% | 21.42% | 71.4% | |||||||||||||
| Accuracy (Non Rejected) | 100% | 91.60% | 54.45% | 0% |
We measured overall accuracy, rejection rate, and accuracy among non-rejected responses. Overall accuracy is defined as the model’s ability to correctly differentiate between real and fake news headlines and their associated images across the four conditions. Rejection rate is defined as the frequency at which the model declines to provide an answer for a given prompt. Non-rejected accuracy is defined as the model’s accuracy in cases where it provided a response. The combined prompt with Google search for headlines and image analysis for images passed to GPT-4o (Condition 1) achieved optimal performance with 100% accuracy and 0% rejection rate on the sample, leading to its selection for the system design.
3.4. Interaction Flow & Conversation Data
Each session follows a standardized protocol. Participants first indicate whether they have seen the news item before, then provide an initial authenticity rating using a binary choice (REAL/FAKE) and confidence slider (0-100%). The system then initiates a conversation where the AI attempts to persuade users toward the correct answer if they were wrong (refer to Figure 13 to view an example from one such conversation). If they were correct, the AI response reinforces the user’s belief, and the user can still interact with the AI system (refer to Figure 12 to view an example from one such conversation).
After the dialogue concludes (maximum 9 exchanges), participants provide a final rating for the news item using the same interface. All interactions are logged with timestamps, including initial ratings, conversation turns, and final ratings. The system records participant responses, AI responses, and metadata (participant ID, item details) for subsequent analysis. Data is stored both on Qualtrics and via the Google Sheets API. Appendix Section D provides additional examples and details on the conversation data.
4. Study Design
We conducted a longitudinal study to understand how AI assistance affects people’s ability to detect misinformation over time. The study ran for four weeks with three testing sessions at weeks 0, 2, and 4. Each session had three phases: (i) Before AI: participants assessed news items independently, (ii) With AI: they interacted with our AI chatbot about the same news items, and (iii) After AI: they independently evaluated new unseen items without AI assistance. This design enabled us to measure both immediate AI assistance effects and lasting skill development—distinguishing between momentary performance gains and genuine improvement in their independent detection abilities. Figure 5 shows the experimental setup.
4.1. Participants
We recruited 105 participants through Prolific, an online research platform, targeting English-proficient adults aged 18 and older from the US and UK. Of the initial 105 participants, 28 dropped out before completing all three sessions, and we excluded an additional 10 participants who failed more than one attention check per session. Attention checks were implemented to ensure participant engagement throughout the study. This resulted in a final sample of 67 participants who completed the entire longitudinal study and provided reliable data for analysis.
The final sample had a mean age of 40 years (range: 20-63) with relatively balanced gender representation (53.7% male, 46.3% female). Participants were predominantly from the United States (76.1%) and the United Kingdom (23.9%). The ethnic composition included White participants (56.7%), Mixed ethnicity (14.9%), Black participants (14.9%), Asian participants (6.0%), Other ethnicities (4.5%), and Unknown ethnicity (3.0%). Figure 11 presented in the Appendix section C displays the complete demographic distribution of the study sample.
Participants received $8 compensation for each 40-minute session and a $3 bonus for completing all three phases. This study was approved by our Institutional Review Board at MIT (Protocol #2505001647), and all participants provided informed consent.
4.2. Experimental Protocol
Each weekly session followed the same three-phase protocol. In the Before AI phase, participants provided initial authenticity ratings on 4 randomly selected headline-image pairs (2 real, 2 fake) without any AI assistance. Next, in the With AI phase, participants interacted with the AI chatbot on the same 4 headline-image pairs for up to 9 rounds, engaging in dialogue to refine their judgments and potentially change their initial assessments. Finally, in the After AI phase, participants evaluated previously unseen items without AI assistance to measure independent skill development. In week 0, participants evaluated 4 new items, while in weeks 2 and 4, they evaluated 6 items: 4 new unseen items and 2 items from the previous week to assess retention. This design allowed us to capture baseline performance, immediate AI assistance effects, and whether participants developed skills for misinformation detection.
4.3. Data Collection
We collected quantitative, qualitative, and exploratory data to assess both performance and participant characteristics. Our primary measures included accuracy, scored as binary classification performance (1 = correct, 0 = incorrect), confidence measured as self-reported certainty on a 0-100% scale, and belief change calculated as the difference between initial and final ratings during AI assistance.
We collected additional measures to characterize our sample and verify data reliability for every participant. For attention checks, we implemented a two-stage process: first, participants who failed 2 pre-interaction attention checks were removed from the study, and second, after each AI interaction, participants answered 2 additional attention checks, leading us to exclude 10 participants who failed these post-interaction checks. We measured participants’ prior AI experience using three questionnaires: (i) Frequency of AI chatbot usage (ChatGPT, Gemini, Claude, Perplexity) from ”Never” to ”More than 6 hours a day,” (ii) Frequency of AI image generation tool usage (ChatGPT, Midjourney, Meta AI) from ”Never” to ”Everyday,” and (iii) AI literacy that contains questions like ”I understand the basic concepts of AI” and rated from strongly disagree to strongly agree. We present additional details in Appendix E Figure 15. We also collected qualitative feedback through three open-ended questions about whether participants felt the AI taught them detection skills, their thinking process during AI dialogue, and any technical or other issues they experienced.
5. Analysis
Our dependent variable was item-level accuracy (, ). Analyses were restricted to participants who completed all three phases (before, with, after) at Weeks 0, 2, and 4 and did not fail more than one attention check (n=67). Items flagged as ”seen before” were excluded from analysis.
We estimated treatment effects using linear probability models (Ordinary Least Squares regression on accuracy) with participant fixed effects and Heteroskedasticity-Consistent 2 (HC2) robust standard errors, following prior work on AI dialogues and belief updating (Costello et al., 2024). This approach yields interpretable percentage-point effects while absorbing stable individual differences in baseline accuracy.
Our main tests focus on two planned within-week contrasts: (i) , the in-session gain on assisted items when participants conversed with the AI, and (ii) , the unaided learning to new items following interaction. These contrasts are estimated separately for each week and reported in percentage points with HC2 robust 95% confidence intervals.
To capture changes in unaided performance over time, we report our results on after-phase trials only, testing absolute differences in unaided accuracy across weeks (Week 4 vs. Week 0; linear trend). This is done through a linear mixed-effects regression model to assess changes in unassisted accuracy over time, focusing specifically on the “after AI support”. The model includes week as a continuous predictor (coded as 0, 2, 4 for the three time points) and participant fixed effects to control for individual differences in baseline accuracy. We use heteroskedasticity-consistent (HC2) robust standard errors to account for potential variance heterogeneity across participants and time points. The linear trend coefficient, , represents the change in accuracy per two-week interval, allowing us to quantify the rate of decay in unassisted performance. We additionally test the specific contrast between Week 4 and Week 0 using a linear hypothesis test with the same robust covariance matrix to confirm the overall decline from initial to final time points. For descriptive context, we plot unaided accuracy by week with Wilson 95% confidence intervals.Because skill development may differ by veracity, we replicated the after-only analyses separately for fake versus real items. All tests are two-sided with .
To identify which conversational strategies influenced fake news detection performance, we applied LLM-as-a-judge methodology to classify each Human-AI dialogue into 21 binary strategy categories. These categories captured specific conversational moves across six domains: general behaviors (e.g., asking guiding questions, probing deeper), evidence strategies (e.g., image forensics, source checking), reasoning strategies (e.g., logic plausibility checks, alternative explanations), emotional strategies (e.g., bias awareness, emotional trigger detection), knowledge activation (e.g., prior knowledge recall, fact comparison), and metacognitive strategies (e.g., confidence calibration, process reflection). Each classifier used structured prompts with explicit descriptions, examples, and binary classification options (0/1) (See all prompts in Appendix E.2). We applied reliability filtering to ensure robust correlations, including only strategies with occurrence rates between 5% and 95%, at least 10 positive classifications, and sufficient variance (¿0.05).
We calculated Pearson correlation coefficients, , between each conversational strategy (binary coded by LLM-as-a-judge) and accuracy outcomes under two conditions: AI-assisted accuracy (during Human-AI interaction) and unassisted accuracy (after AI interaction). Correlations were computed separately for fake news items to focus on misinformation detection. For each qualifying strategy, we calculated the difference between unassisted and AI-assisted correlations (diff = - ) to identify whether strategies promoted learning transfer (positive difference) or AI dependence (negative difference). We also computed standard deviations from the strategy occurrence distributions to assess variability across conversations.
6. Quantitative Results
6.1. Descriptive Overview
The final sample consisted of 67 participants who completed all three weekly sessions (Weeks 0, 2, and 4), from the initial 105 recruited participants. The analysis included 4,536 total item ratings across the three phases (1,512 ratings per phase). Participants engaged in an average of 9 conversational turns with the AI system per news item, generating 7,203 pairs of Human-AI conversations. Additional details on conversations are present in Appendix section D.
6.2. Persuasive AI substantially improved accuracy when assisting the user
In our results, participants showed large short-term gains when interacting with the persuasive AI. Across all weeks, accuracy on assisted items was markedly higher than before assistance, with an average uplift of about percentage points (pp) (Week 0: , Week 2: pp, Week 4: pp). These week-by-week effects were robust, and the assisted gain was somewhat larger at Week 4 compared to Week 0 ( pp). This pattern suggests that the AI chatbot consistently corrected misperceptions during conversations and that the strength of this persuasion effect may have grown over time. See Figure 6.
6.3. People’s accuracy with AI did not persist when they were unassisted, and their detection abilities declined
Our results show that after interacting with the AI when it is unavailable, their accuracy on unaided items (”After AI support”) declined steadily over the four-week period. When judging new items without AI assistance, accuracy decreased linearly across weeks (, , , ). A direct comparison confirmed that Week 4 accuracy was significantly lower than Week 0 (, ). These results suggest that repeated exposure to AI support improved participants’ judgments with the AI chatbot but eroded their ability to independently identify misinformation when they no longer had access to it.
6.4. Declines in abilities were driven by failures to detect fake content
The decline in accuracy when unassisted was driven almost entirely by performance on fabricated news: Accuracy on fabricated items fell sharply over time (, , , ), while accuracy on real news items remained stable (, , , ). This suggests that the deterioration in unaided performance was driven specifically by a reduced ability to identify false content, while recognition of real news did not change.
7. AI Interaction Strategies and Fake News Discernment
To understand which conversational strategies shaped fake news detection, we applied LLM-as-a-judge (Gu et al., 2024; Zheng et al., 2023) to classify each of the human–AI dialogues into 21 classifier categories. These classifiers capture specific conversational moves such as asking for image forensics, probing deeper, or prompting confidence calibration (see Appendix E.2 for full descriptions). For example, image forensics questions occur when the AI asks about visual artifacts or manipulations in an image (e.g., shadows or lighting), while confidence calibration refers to prompting the user to rate and reflect on how certain they are in their judgment. By coding conversations in this way, we were able to correlate strategy use with accuracy on fake items under both assisted and unassisted conditions. To avoid correlations with rarely occurring strategies, our analysis included only strategies with occurrence rates between 5% and 95% of all conversations with at least 10 positive classifications, and a sufficient variance (). We then calculated Pearson correlation (r) coefficients between each conversational strategy (binary coded by LLM-as-a-judge) and accuracy outcomes under two conditions: AI-assisted accuracy (during Human-AI interaction) and unassisted accuracy (after AI interaction). For each qualifying strategy, we computed correlations separately for fake news items and calculated the difference between unassisted and AI-assisted correlations (diff = - ) to identify whether strategies promoted learning transfer (positive difference) or AI dependence (negative difference). Standard deviations were calculated from the strategy occurrence distributions to assess variability across conversations. An overview of these correlations can be found in Figure 7.
7.1. Strategies paying attention to image artifacts and prior knowledge were associated with higher assisted accuracy
During AI-assisted sessions, strategies that directed attention to concrete cues were the strongest predictors of success. Image forensics questions (, occurring in 22.9% of conversations, ) and recall of prior knowledge (, 68.8%, ) improved assisted accuracy the most. More modest gains came from logic plausibility checks (, 37.5%, ), source checking prompts (, 18.8%, ), and contextual consistency checks (, 33.3%, ). Other strategies had negligible effects (reflect on process, , 43.8%, ) or negative correlations, including confidence calibration (, 43.8%, ), probing deeper across multiple messages (, 91.7%, ), and guiding questions (, 83.3%, ). In short, when the AI helped users look for manipulation or connect back to what they already knew, participants were better able to identify fake content.
7.2. Guiding and probing questions were associated with independent detection abilities
On unaided trials after using the AI, a different pattern emerged. Guiding questions—where the AI in previous interactions had asked users previously to reflect on sources or verifiable details—flipped from negative in assisted sessions () to the strongest predictor of unassisted detection (; diff=, std=, 83.3%, ). Similarly, probing deeper over multiple messages was associated with reduced accuracy while interacting with the AI () but was associated with higher accuracy after interacting with the AI unassisted (; diff=, std=, 91.7%, ). Forensics and knowledge cues again carried over, with questions around image forensics during AI interaction increased both accuracy with the AI and when unassisted afterwards ( with AI vs. unaided; diff=, std=, ). The AI asking the user to recall prior knowledge that is relevant also increased accuracy with and unassisted after interacting with the AI ( with AI vs. unaided; diff=, std=, ).
7.3. Some strategies helped in the moment but undermined learning
Several strategies proved to be only helpful when the AI chatbot was present. Confidence calibration is the clearest case: it slightly reduced accuracy with AI (, 43.8%, ) and strongly undermined unassisted accuracy after AI (, diff=). Similarly, devil’s advocate roleplay provided no assisted benefit (, 12.5%, ) and worsened unaided performance (). These strategies appear to foster dependency by drawing attention to subjective confidence or rhetorical sparring rather than transferable heuristics.
7.4. Balanced strategies supported both assistance and learning
Finally, some conversational moves provided small but consistent benefits across different news items. Image forensics questions (with AI , unassisted after AI , diff=, std=, ) and recall of prior knowledge (with AI , after AI , diff=, std=, ) boosted accuracy regardless of whether the AI was present. Similarly, logic plausibility checks (with AI , after AI , diff=, ) and ask-for-source checking (with AI , without AI , diff=, ) showed modest positive effects in both conditions. These “balanced strategies” offered steady but modest returns, reinforcing their value as design levers for both immediate support and independent discernment.
8. Keyword Extraction Based Conversation Analysis
To systematically identify conversational patterns between subjects and AI, we employed keyword matching techniques commonly used in computational linguistics (Manning and Schutze, 1999; Keselj, 2009). Drawing upon prior works (Maragheh et al., 2023), we automated the keyword extraction process by providing Claude 3.5 Sonnet with conversation transcripts and prompting it to generate comprehensive keyword lists for three key behavioral dimensions: (i) agreement patterns (human agreeing with AI versus AI agreeing with human), (ii) independent thinking (human versus AI independent reasoning), and (iii) probing behavior (human questioning AI versus AI questioning human). These keyword lists were then applied to all conversation transcripts to quantify behavioral patterns throughout the study period.
8.1. Categories and Examples
Our analysis identified three primary behavioral categories that characterize Human-AI conversational dynamics. These categories could help explain potential mechanisms underlying the performance degradation of misinformation discernment accuracy over time: increased agreement with the system over time, reduced questioning behavior, and decreased independent thinking. Agreement behavior captures instances where one party validates the other’s perspective, distinguishing between human agreement with AI versus AI agreement with human responses. People typically use casual affirmations (“you’re right,” “that makes sense,” “totally”) while AI employs more formal validation (“you make a valid point,” “your perspective is valid,” “I agree with your assessment”). Independent thinking encompasses expressions of personal reasoning and uncertainty. People demonstrate subjective stance-taking (“I think,” “I believe,” “maybe”) compared to AI’s analytical hedging (“it seems,” “this suggests,” “considering”). Probing behavior reflects information-seeking and curiosity. People ask direct, challenging questions (“why do you,” “how do you know,” “are you sure”) while AI engages in Socratic inquiry (“what do you think,” “help me understand,” “have you considered”). We include the complete keyword list in the Appendix section E.3 Figure 16, 17, 18.
8.2. Insights
Figure 8 shows the longitudinal analysis of conversations, revealing shifts in conversational dynamics across the month-long period. Human agreement with AI increased from 23.0% in Week 0 to 28.5% in Week 4, while AI agreement with humans remained relatively stable (4.2% to 6.4%), suggesting growing human acceptance of AI perspectives.
Conversely, human independent thinking remained consistently low across all weeks (7.2% to 7.6%), while AI independent thinking increased from 19.8% to 25.1%, indicating that the AI maintained analytical autonomy while humans’ independent thinking ability did not vary much. The probing behavior showed a decline in curiosity over time, with human questioning of AI increasing slightly (from 15.6% to 19.9%) and then declining to 17.6% in Week 4, while AI probing of humans remained relatively stable (25.6% to 24.6%).
These conversational patterns align with our broader findings that AI assistance provided immediate performance benefits during AI-assisted sessions, but participants’ independent reasoning abilities deteriorated over the month-long study period. Together, this suggests a potential overreliance on AI guidance at the expense of fake news discernment skills for misinformation detection.
9. Qualitative Analysis
A thematic analysis of the conversation transcripts was conducted to find overlapping themes among participants. Inductive coding (Boyatzis, 1998) was conducted to discover salient themes. Two coders iteratively coded the narratives with themes, discussed and aligned any disagreements, and added emerging themes until reaching theoretical saturation (Creswell and Creswell, 2017).
To understand how individual participants’ relationships with AI assistance evolved over time, we employed a longitudinal thematic approach examining within-person change across the 4-week study period. Rather than analyzing aggregate patterns, we tracked individual participant trajectories by comparing their Week 0 and Week 4 responses for changes in: (i) AI Effectiveness: Participants responded to “Do you feel that the AI taught you how to identify true and false information and images? Why? Why not?”, and (ii) Thinking Process: Participants responded to “Please describe your thinking process when reacting to the AI dialogue? Be as detailed as possible.”
Each participant’s trajectory was characterized by coding their initial and final responses (Week 0 and Week 4) along these two dimensions, then identifying patterns of change and categorizing all 67 participants into five distinct trajectory types. Our analysis revealed the following five distinct trajectory patterns described below:
Progressive Learners (n=18, 27%). These participants began expressing skepticism about the AI assistant but progressively developed confidence in AI. Multiple participants exhibited this pattern (P9, P12, P14, P17, P27).
At week 0, participants expressed skepticism and caution, but in week 4, they started trusting AI. For instance, ”No, I don’t feel it taught me. The AI didn’t cite its sources. It didn’t give me scientifically proven information either.” (P17).At week 4, ”Yes. The AI taught me how to identify true information by cross-referencing many credible sources. If an article is only available on one site and you search and nothing else shows up for that, it’s most likely to be false info.” (P17).
Consistently Collaborative Learners (n=12, 18%).These participants maintained positive engagement with AI throughout all four weeks while developing systematic approaches to incorporating AI feedback into their decision-making. However, their consistent collaboration showed signs of developing dependence on AI validation.
For example, Participants in this group (P21, P28, P32, P37) consistently tried learning from AI and demonstrated collaborative behavior.
At Week 0, ”I think it reinforced my precautions I already try to use” (P21),
”Yes, I think it taught me how to identify true and false information” (P32).By week 4, they reinforced similar comments as week 0. ”It helped remind me to be skeptical and logical”(P21),
”Yes, it taught me to first confirm sources before I make my decision as to what is real or fake” (P32).While maintaining positive learning attitudes, the evolution toward increased reliance on AI-provided tips and sources suggests a growing dependence on AI for validation rather than the development of independent judgment.
Growing Skeptics (n=8, 12%). These participants began with openness to AI assistance but, with time, became more critical and explicitly asserted their independence from AI guidance. This trajectory is characterized by increasingly dismissive language about AI’s contribution to their abilities.
At Week 0, these participants (P2, P6, P23, P48) showed initial engagement:”I think the AI gave me some helpful advice. It told me to really focus on what the headline is actually saying” (P2),
”Yes, I feel that they gave me some useful ideas to help me make these determinations” (P48).By Week 4, their stance had shifted markedly: ”I feel like the AI just gave me suggestions based on the information from its sources. It didn’t really teach me how to differentiate between true and false information” (P2), ”They didn’t teach me much about exploring the context of the images themselves…I was fairly passive in this process, while the AI chatbot fact-checked the stories” (P48).
This reassertion of cognitive autonomy and explicit recognition of passivity may indicate these participants recognized their declining independent performance when engaging with AI and compensated by distancing themselves from AI influence.
Persistent Self-Reliant Learners (n=15, 22%). These participants consistently maintained their independence throughout the study, explicitly positioning AI as a secondary tool for confirmation rather than a primary source of guidance.
For instance, at Week 0, these participants (P6, P7, P51) established clear boundaries with statements emphasizing personal agency:
”My ability to determine true/false info comes from before AI was popularised—it’s based on my own intuition and critical thinking skills” (P6),
”When I was engaging with the AI, my main goal was to double-check and confirm the choices I was already leaning toward. I didn’t just rely on the AI to give me the answer; instead, I used it as a tool to validate or challenge my own reasoning” (P51).By Week 4, ” I feel like I can tell quite quickly if most images are AI generated or not, and can notice most fake news based on the fact I read the news frequently from a variety of sources.” (P6),
”When reacting to the AI dialogue, I first read the claim carefully to understand the who, what, where, and when…I compare the claim with what I know from credible sources or memory”(P51).While this independence preserved critical thinking skills, the consistent resistance to AI guidance may have prevented these participants from benefiting from legitimate learning opportunities.
Dependency Developers (n=14, 21%). These participants demonstrated progressive shifts from active self-reliance to passive acceptance of AI guidance, with language increasingly indicating cognitive offloading and reduced personal effort in decision-making.
At Week 0, participants like P48 showed active engagement: ”Yes, I feel that they gave me some useful ideas to help me make these determinations. I still am not perfect at this, but I think I am improving.”
By Week 4, their descriptions revealed concerning passivity: ”They didn’t teach me much about exploring the context of the images themselves. They did emphasize that you must check across multiple sources to make sure a story is true. They did this for me, so I was fairly passive in this process, while the AI chatbot fact-checked the stories. ”
The explicit acknowledgment of being ”fairly passive” represents the concerning development of cognitive dependency, where participants shifted from active learning to passive acceptance of AI-performed verification.
10. Discussion
10.1. Changing beliefs about misinformation vs. improving abilities to detect misinformation
Our findings reveal a critical distinction between correcting false beliefs and building lasting detection capabilities. Previous work by Costello et al. (Costello et al., 2024) demonstrated that AI dialogues can reduce belief in conspiracy theories by 22%. Our results confirm that AI can effectively change beliefs about specific misinformation items with participants showing an average of 21% improvement in accuracy during AI-assisted sessions across all weeks. However, this immediate belief correction did not translate into improved independent discernment abilities. While participants benefited from AI’s persuasive reasoning (+21 pp accuracy gain), they later became less able to evaluate new misinformation on their own (–15.3 pp decline by week 4). Interestingly, this decline was concentrated on detecting fake content, suggesting that once participants learned to trust AI for verification, they lost vigilance in scrutinizing suspicious claims themselves. This has implications for how we currently use AI to assist us in information processing. If the goal is simply to reduce the belief in specific falsehoods “in the moment,” AI can be effective. But if the goal is to build people’s skills to recognize falsehoods on their own, our results suggest current AI approaches may fail—or even backfire by fostering reliance.
10.2. Short-term gains vs. long-term learning
The temporal dynamics of our findings align with broader research on Human-AI collaboration, showing immediate performance benefits that mask underlying skill degradation. Similar patterns have been observed in other domains where AI assistance provides cognitive scaffolding — users often experience enhanced performance during assisted trials but then develop dependency that undermines their capabilities to think on their own (Danry et al., 2023; Agarwal et al., 2025; Kosmyna et al., 2025). Similarly, a recent study also found that doctors using AI classification tools to detect cancer became worse at detecting cancer alone after having used the tool (Budzyń et al., 2025).
10.3. Skill deterioration and over-reliance
The results demonstrate that participants readily incorporated AI assistance into their decision-making process but failed to internalize the underlying detection strategies. Instead of developing independent discernment skills, participants increasingly relied on AI validation, as evidenced by both quantitative performance data and qualitative trajectory analysis. The 27% of participants classified as ”Progressive Learners” (See Section 9) exemplify this concerning pattern — they self-reported enhanced detection abilities while experiencing actual performance decline as measured in quantitative analysis. This skill deterioration was particularly pronounced for detecting fake content, however, for real content detection, it remained stable, suggesting that AI assistance specifically undermined participants’ ability to identify sophisticated misinformation rather than affecting general news comprehension. This pattern indicates that the AI’s persuasive approach may have created false confidence about fake news detection without building the critical thinking skills necessary for independent evaluation. Our conversational analysis provides mechanistic insight into this dependency development. The progression from active questioning (15.6% to 19.9% human probing in early weeks) to passive agreement (23.0% to 28.5% human agreement by Week 4) illustrates how participants gradually shifted from collaborative inquiry to uncritical acceptance. Meanwhile, participants’ independent thinking remained consistently low (around 7%), indicating that the AI’s persuasive approach failed to stimulate autonomous reasoning processes.
10.4. Towards an AI chatbot that improves discernment abilities
Our findings highlight the need for AI systems designed specifically to enhance rather than replace human reasoning capabilities. Recent work suggests that Socratic questioning methods may better support skill development by encouraging active learning rather than passive acceptance of information (Danry et al., 2023).
The trajectory analysis reveals that some participants (the 12% classified as ”Growing Skeptics”) successfully maintained critical distance from AI assistance while still benefiting from the interaction. Understanding the characteristics and strategies of these participants could inform the design of AI systems that promote healthy Human-AI collaboration in misinformation detection contexts.
Additionally, our findings suggest the need for hybrid approaches that combine immediate belief correction with long-term skill building. While persuasive AI effectively addresses false beliefs in the moment, complementary interventions focused on critical thinking skills may be necessary to build lasting discernment capabilities. The implications extend beyond misinformation detection to any domain where AI assistance aims to augment human judgment. Our methodology for distinguishing between AI-assisted performance and independent skill retention provides a framework for evaluating whether AI tools genuinely enhance human capabilities or merely create performance dependencies. As AI systems become more sophisticated and persuasive, ensuring they build rather than undermine human discernment becomes increasingly critical for maintaining informed democratic discourse.
11. Limitations and Future Work
Dataset Size. Beginning with 55 randomly selected items from MiRAGeNews, we removed 6 items (10.9%) where our AI system failed to produce correct classifications, leaving only 49 validated news items for the main study. This relatively small dataset may limit the generalizability of our findings across different types of misinformation and news domains.
Participants. Our participants were recruited from Prolific and were predominantly from the US and UK, potentially limiting generalizability to other populations, cultures, and languages.
Study Duration. While our month-long study captured important longitudinal trends, longer-term studies would be valuable to understand whether the observed dependency effects could persist, worsen, or eventually reverse after extended periods without AI assistance.
Several promising avenues emerge from our limitations. Longitudinal studies extending beyond four weeks could reveal whether skill degradation continues or whether recovery occurs after AI assistance withdrawal. Cross-cultural validation would establish whether dependency effects vary across different educational systems, media landscapes, or cultural approaches to authority and expertise. Comparative studies examining different AI pedagogical approaches — particularly Socratic questioning methods that prompt user reasoning rather than providing direct answers — could identify design principles that enhance rather than undermine skill development.
12. Conclusion
This study reveals a concerning paradox in AI-assisted misinformation detection: while persuasive AI systems produce immediate accuracy gains (+21%), they simultaneously undermine long-term independent judgment (-15.3% by week 4). Our longitudinal analysis demonstrates that current approaches prioritize belief correction over skill development, creating dependency rather than durable discernment capabilities. The findings challenge assumptions about AI’s educational potential and highlight the need for systems designed to enhance rather than replace human reasoning. As AI becomes increasingly sophisticated, ensuring these tools build critical thinking skills rather than cognitive dependency becomes essential for maintaining public resilience to misinformation.
Appendix A Misinformation Spread and Impact
In May 2023, AI-generated reports and a fake image of an explosion near the Pentagon spread rapidly on social media, garnering thousands of shares before being debunked. The realistic-looking doctored photograph showed smoke rising near the Pentagon, causing brief public panic and temporary stock market fluctuations as Dow futures dropped momentarily (See Figure 9). Pentagon officials and authorities quickly confirmed no incident had occurred. The episode demonstrated how sophisticated AI-generated misinformation can appear credible enough to fool users and trigger real-world economic consequences, highlighting urgent concerns about AI’s potential for creating and rapidly spreading false information on an unprecedented scale.
Appendix B Prompts
Figure 10 describes the three prompts used during the system design experiment.
Appendix C Participants
Figure 11 presents the demographic information about the participants.
Appendix D Conversation Details
D.1. Conversation Volume and Length
Our study generated a total of 7,203 Human-AI conversation pairs across the three experimental sessions. The distribution of conversations by week was as follows:
Week 0: 2,262 conversations
Week 2: 2,466 conversations
Week 4: 2,475 conversations
The average number of conversational turns averaged out to 9 across the weeks, indicating consistent engagement patterns throughout the longitudinal study.
D.2. Conversation Characteristics
All conversations were automatically logged with timestamps, participant responses, AI response, and associated metadata. Each conversation was structured around a specific news item, with participants initially providing their assessment and the AI system responding with evidence-based persuasive dialogue tailored to their reasoning.
The conversations typically followed a pattern where participants would state their initial belief about a news item’s authenticity, the AI would probe their reasoning, provide evidence or counterevidence, and engage in iterative dialogue aimed at improving the participant’s detection accuracy. The relatively consistent conversation length across weeks indicates stable engagement patterns throughout the longitudinal study. Some of the examples of these conversations are present in Figure 12, 13 and 14.
Appendix E Quantitative Analysis
E.1. AI usage and AI Literacy versus Accuracy Graph
We assessed participants’ AI literacy using a six-item scale with 7-point Likert responses ranging from ”Strongly Disagree” to ”Strongly Agree.” The scale measured participants’ self-reported understanding and confidence with AI concepts across multiple dimensions. Items included: basic AI concept comprehension (”I understand the basic concepts of AI”), self-efficacy in AI project contribution (”I believe I can contribute to AI projects”), critical evaluation skills (”I can judge the pros and cons of AI”), knowledge currency (”I keep up with the latest AI trends”), social comfort with AI discussions (”I’m comfortable talking about AI with others”), and creative application abilities (”I can think of new ways to use existing AI tools”). This multi-faceted approach captured both technical understanding and practical confidence with AI technologies.
We also measured participants’ AI usage frequency across two domains using ordinal scales. For AI chatbot usage (ChatGPT, Gemini, Claude, Perplexity), participants selected from six frequency categories ranging from ”Never” to ”More than 6 hours a day,” with intermediate options including ”Once a month,” ”Once a week,” ”0-2 hours a day,” ”2-4 hours a day,” and ”4-6 hours a day.” For AI image generation tool usage (ChatGPT, Midjourney, Meta AI), participants chose from five frequency categories: ”Never,” ”Rarely,” ”Once a month,” ”Once a week,” and ”Everyday.” These measures captured both text-based and visual AI tool engagement, providing insight into participants’ practical experience with different AI modalities that could influence their fake news detection abilities. We present the findings in Figure E.1.
E.2. LLM-as-a-judge Classifier Definitions
For analysis, we applied a set of LLM-as-a-judge classifiers to label conversational strategies in each human–AI dialogue. Below we summarize the categories and provide definitions and examples of each strategy.
E.2.1. General Behaviors
Gave away ground truth: Whether the AI explicitly revealed the correct answer (e.g., “This headline is false/true because…”).
Asked broad questions: Open-ended prompts such as “What do you think about this headline?”
Asked guiding questions: Targeted prompts to steer reasoning (e.g., “What details can you verify?”).
Probed deeper: Follow-up questions building on user responses (e.g., “Can you tell me more about why you think that?”).
Changed focus: Switching topics each turn without continuity.
E.2.2. Evidence Strategies
Ask for source checking: Prompting users to consider the credibility of the source.
Cross-verification prompting: Encouraging users to verify information via other outlets.
Image forensics questions: Asking about visual manipulations (e.g., lighting, shadows).
Contextual consistency checks: Comparing whether the image matches the headline context.
E.2.3. Reasoning Strategies
Counterfactual generation: Asking users to imagine different scenarios if the headline were true vs. false.
Alternative explanations: Encouraging users to propose more mundane explanations.
Logic plausibility checks: Asking whether claims make sense given known facts.
Scale/scope reflection: Asking about the likely impact if the claim were true.
E.2.4. Emotional Strategies
Spot emotional triggers: Asking if the content is designed to provoke strong emotions.
Bias awareness reflection: Highlighting potential confirmation bias.
Framing analysis: Asking how the headline is framed (e.g., sensational, vague).
E.2.5. Knowledge Activation
Recall prior knowledge: Prompting recall of what the user already knows.
Fact comparison: Comparing claims against known facts or statistics.
Historical analogies: Asking whether similar events have occurred before.
E.2.6. Questioning Types
Devil’s advocate roleplay: Presenting an opposite stance for the user to respond to.
Evidence ranking: Asking users to rate the strength of different pieces of evidence.
Step-by-step reasoning: Guiding users to articulate reasoning steps sequentially.
E.2.7. Metacognitive Prompts
Confidence calibration: Asking users to rate their confidence (e.g., “On a scale of 1–10, how sure are you?”).
Uncertainty tolerance: Normalizing that uncertainty and verification are acceptable.
Reflect on process: Asking users to review how they arrived at a decision.
E.3. Keywords for Agreeing, Probing, and Independent Thinking